OpenStudy (anonymous):

sqrt x^2-5x+4=x-3 solve the radical equation

OpenStudy (anonymous):

\[\sqrt(x^2-5x+4)=x-3 ?\]

OpenStudy (anonymous):

Is it this you mean? \[\sqrt{x ^{2}-5x+4}=x-3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Just square both sides then

OpenStudy (anonymous):

square both sides and get: x^2-5x+4=(x-3)^2 x^2-5x+4=x^2-6x+9 x=5